OpenStudy (anonymous):

Diiferentiate \[f(x)=(2x^5-4x^4+3)^{308}\]

OpenStudy (anonymous):

ok so would it be \[f'(x)=308(2x^5-4x^4+3)^{307}\] or do you do that and ALSO take the derivative within the parenthesis like \[f'(x)=308(10x^4-16x^3)^{307}\]

sam (.sam.):

Yes both includes, but there's a bit error in your calculations, it should be \[f'(x)=308(10x^4-16x^3)(2x^5-4x^4+3)^{307}\]

OpenStudy (anonymous):

ok thanks